# Social Bias in Large Language Models For Bangla: An Empirical Study on Gender and Religious Bias

### v2

@@ Abstract
The rapid growth of Large Language Models (LLMs) has put forward the study of biases as a crucial field. It is important to assess the influence of different types of biases embedded in LLMs to ensure fair use in sensitive fields. Although there have been extensive works on bias assessment in English, such efforts are rare and scarce for a major language like Bangla. In this work, we examine two types of social biases in LLM generated outputs for Bangla language. Our main contributions in this work are: (1) bias studies on two different social biases for Bangla (2) a curated dataset for bias measurement benchmarking (3) testing two different probing techniques for bias detection in the context of Bangla. This is the first work of such kind involving bias assessment of LLMs for Bangla to the best of our knowledge. All our code and resources are publicly available for the progress of bias related research in Bangla NLP.  [^1] https://github.com/csebuetnlp/BanglaSocialBias
@@ 1 Introduction
Figure 1: Workflow for the creation of corpus for the experiments detailed in this study.
@@ 2 Related Work
Existence of gender bias has been exposed in tasks like Natural Language Understanding Bolukbasi et al. (2016); Gupta et al. (2022); Stanczak and Augenstein (2021) and Natural Language Generation Sheng et al. (2019); Lucy and Bamman (2021); Huang et al. (2021). Benchmarks such as WinoBias Zhao et al. (2018) and Winogender Rudinger et al. (2018) have been used to measure gender biases in LMs. Preliminary studies on religious and ethnic biases are done in some works (Milios and BehnamGhader, 2022; Navigli et al., 2023; Abid et al., 2021). Works like Nadeem et al. (2021); Nangia et al. (2020) provide frameworks and datasets for different types of biases including gender and religion. IndiBias Sahoo et al. (2024), a benchmark in Indian context, has been introduced to measure socio-cultural biases in LLMs.
Recent studies have conducted experiments on determining gender stereotypes in LLMs Kotek et al. (2023); Ranaldi et al. (2024); Jha et al. (2023); Dong et al. (2024a) and debiasing techniques Gallegos et al. (2024); Ranaldi et al. (2024), but most of them are on English. There are a few works on multilingual settings Zhao et al. (2024); Vashishtha et al. (2023), but such efforts are not common for Bangla. The most preliminary work on Bangla bias detection is found in the works of Sadhu et al. (2024), that includes static and contextualized embeddings. Effectiveness of varied probing techniques for extracting cultural variations in PLMs has been discussed in Arora et al. (2023).
@@ 3 Data
We use two strategies for LLM probing: Template Based and Naturally Sourced. The workflow of our proposed approach is illustrated in Figure 1. We explain the two techniques as follows:
Template Based: We create semantically bleached templates with placeholders for specific traits that are filled with adjective words to create prompts. We use these prompts for gendered role or religious identity prediction. The adjective words encompass different categories used to describe human nature e.g. Personality, Outlook and Communal based. We also include Occupation as a category to understand the association between gender and religion with occupation. We create templates to intermix human professions with traits to study if the addition of occupation affects role prediction. We ensured that no stereotypes were induced into the templates and that all adjective words and occupations are equally probable for any gender or religious community. In total, we have 2772 template sentences by combining both the categories (see appendix C.1 for more details).
Naturally Sourced: We use the BIBED dataset Das et al. (2023), specifically the Explicit Bias Evaluation (EBE) data for naturally occurring scenarios. The sentences are structured in pairs, each containing one identifying subject from a group of either male-female words (for gender) or Hindu-Muslim words (for religion). We replace the main identity dimension in the sentence with _ (gap) and give the model options to select between the possible identities. We provide details of the procedure and examples in appendix C.2. However, there are many instances where a sentence is not equally probable for both contrasting identities. In order to curate sentences that serve our cause, which is mainly to provide equal opportunity, we manually select these sentences (refer to appendix D for details). After the curation process, we are left with 2416 pairs for gender and 1535 pairs for religion.
@@ 4.1 Model Selection
For our experiment we provide results for four state-of-the-art LLMs: Llama3-8b (version: Meta-Llama-3-8B-Instruct 333meta-llama/Meta-Llama-3-8B-Instruct) (AI@Meta, 2024), GPT-3.5-Turbo 444gpt-3-5-turbo, GPT-4o 555gpt-4o and Claude-3.5-Sonnet666anthropic/claude-3.5-sonnet. Since Bangla is a low resource language, not many models could generate the expected response we required. We also experimented with some other models (mentioned in limitations) but none could produce any presentable result serving our purpose.
@@ 4.2 Prompt
We design prompts to best fit the generation task. In the case of template based probing, we prompt the model for gendered role or religious identity selection, and in the case of naturally sourced probing, we use fill in the blanks approach. We provide examples of prompt creation in appendix G. In table 1, we provide the number of unique prompts for each model.
@@
During evaluation, the options provided to LLMs inside a prompt are randomly shuffled for both gender and religious entities to avoid selection bias Zheng et al. (2024).
4.3 Evaluation Setup
Evaluation Metric: We employ the widely used fairness metric, Disparate Impact (DI) Feldman et al. (2015), calculated as $\frac{P(Y=1|S\neq 1)}{P(Y=1|S=1)}$. For our binary identifiers (e.g., male-female, Hindu-Muslim), DI can be applied through empirical estimation. In task Q, for category a with outcomes x and y, DI is calculated as:
@@ 4.3 Evaluation Setup
We use occurrence frequency instead of probability and adjust the metric to adjust equal proportionality in bias scores (further justification and details is provided in appendix B):
@@ 5 Results and Evaluation
Figure 2: Bias Scores in role selection for multiple LLMs in the case of template based probing for gender and religion data. Positive and negative traits results are shown separately. The neutral line $(Bias\>Score=0)$ is highlighted in all the figures. (Note that the results for Occupation are the same for positive and negative traits and only included in contrasting graphs for the ease of comprehending the effect of inter-mixing with other traits.)
@@ 5.1 Template Based Probing Results
Gender Bias: Our findings (Figure 2(a), 2(b)) show that GPT-3.5-Turbo is consistently biased toward females, while Llama-3 and Claude-3.5-Sonnet are biased toward males across both positive and negative traits. GPT-4o exhibits the most fluctuation, switching its bias depending on the category. When the traits change from positive to negative, GPT-4o changes substantially from female direction to male direction for Personality and Communal based traits. Except for GPT-3.5-Turbo, all models display a strong male bias for occupations.
@@ 5.2 Naturally Sourced Probing Results
Figure 3: Bias results in Naturally Sourced(EBE) probing method for multiple LLMs. The upper bound is set to 2
@@
We hypothesize that, the reason for not showing substantial bias in naturally probed examples can be attributed to two points: (1) When Bangla prompt with bigger context is used from a naturally set scenario, the focus of the LLMs are not pin pointed towards the personification of a character into a religious or gender entity. (2) The guard-rails used in LLMs work better in a natural probing setting.
@@ 6 Conclusion
To summarize, by conducting experiments using two different probing techniques and dataset, we investigate gender and religion bias in multilingual LLMs in the context of Bangla. Our work demonstrates the existence of bias in both categories on different degrees. This emphasizes the need for de-biasing techniques to be applied for the use of LLMs in sensitive tasks in the realm of Bangla Language and developing proper linguistics and culturally aware framework for bias measurement. In future, we plan to investigate the effects of bias in downstream applications of Bangla language models.
@@ Appendix B Evaluation Metric Justification
Appendix C Data Preparation
C.1 Preparing Templates
Figure 9 categorizes templates with various placeholder types that we used for gender and religion role prediction. The placeholders are designed to fit different adjective words (illustrated in Figure 6), helping to identify the implied group (gender or religion) through various descriptive angles. We construct curated sentences using the templates and putting in appropriate adjectives and traits. The examples provided in Placeholder column in Figure 9 can be consulted for better understanding.
C.2 Naturally Sourced Sentences
Figure 7 groups sentences into "Gender" and "Religion" biases, providing original (root) sentences, their paired sentences with altered gender or religion, and the necessary modifications to use them as datapoints. For each root sentence, we create a pair sentence by switching gender or religion references. The modifications indicate how placeholders can be used to test for bias, with corresponding answers for each possible substitution.
Appendix D Data Filtration for Naturally Sourced Sentences
The selection criteria for the Explicit Bias Evaluation(EBE) dataset are based on ensuring meaningful and contextually accurate sentences that are neutral from the perspective of gender and religion. In the original BIBED dataset Das et al. (2023), authors created pair for each sentence by replacing the identifying subject, either male-female (for gender) or Hindu-Muslim (for religion) with their respective counterparts (shown in Figure C.2). However, in the EBE data, there are many generated pair sentences that are semantically inconsistent for the pair subject as illustrated in the first two columns of Figure 8.
@@ Appendix D Data Filtration for Naturally Sourced Sentences
Figure 5: Examples of Rejected Sentence and Reason for Rejection
Lastly, we modified each selected root sentence as a data-point by removing its main identifying subject male-female (for gender) or Hindu-Muslim (for religion) and prompted the models for bias detection. The workflow is shown in Figure 5.
Figure 6: Categories of Adjective words used for templates
Figure 8: Examples of Rejected Sentence and Reason for Rejection
Appendix E Annotator’s Agreement on Naturally Selected Data
@@
Appendix F Dataset Statistics
@@
Appendix G Prompt Creation
For each data point, we created prompt template for both template based probing and naturally sourced probing. Each prompt has an instruction followed by an input.
Template Probing: As shown in Table 4, LLMs are instructed to respond with a gender or religion reflecting a Bengali person for template based probing. Each input contains a sentence with gender neutral pronoun along with one of the trait words listed in Figure 6. Input sentence templates with placeholders are explained in Figure 9.
Naturally Sourced Probing: In case for naturally sourced sentences, fill in the blanks approach is taken. LLMs are instructed to fill in the blank with a gender (male-female) or religion (Hindu-Muslim) reflecting the context of the input. Modification of EBE datapoints for prompt creation is shown in Figure 7.

### v3

@@ Abstract
The rapid growth of Large Language Models (LLMs) has put forward the study of biases as a crucial field. It is important to assess the influence of different types of biases embedded in LLMs to ensure fair use in sensitive fields. Although there have been extensive works on bias assessment in English, such efforts are rare and scarce for a major language like Bangla. In this work, we examine two types of social biases in LLM generated outputs for Bangla language. Our main contributions in this work are: (1) bias studies on two different social biases for Bangla, (2) a curated dataset for bias measurement benchmarking and (3) testing two different probing techniques for bias detection in the context of Bangla. This is the first work of such kind involving bias assessment of LLMs for Bangla to the best of our knowledge. All our code and resources are publicly available for the progress of bias related research in Bangla NLP.  [^1] https://github.com/csebuetnlp/BanglaSocialBias
@@ 1 Introduction
Figure 1: Workflow for the creation of naturally sourced corpus for the experiments detailed in this study.
@@ 2 Related Work
Existence of gender bias has been exposed in tasks like Natural Language Understanding Bolukbasi et al. (2016); Gupta et al. (2022); Stanczak and Augenstein (2021) and Natural Language Generation Sheng et al. (2019); Lucy and Bamman (2021); Huang et al. (2021). Benchmarks such as WinoBias Zhao et al. (2018) and Winogender Rudinger et al. (2018) have been used to measure gender biases in LMs. Preliminary studies on religious and ethnic biases are done in some works (BehnamGhader and Milios, 2022; Navigli et al., 2023; Abid et al., 2021). Works like Nadeem et al. (2021); Nangia et al. (2020) provide frameworks and datasets for different types of biases including gender and religion. IndiBias Sahoo et al. (2024), a benchmark in Indian context, has been introduced to measure socio-cultural biases in LLMs.
Recent studies have conducted experiments on determining gender stereotypes in LLMs Kotek et al. (2023); Ranaldi et al. (2024); Jha et al. (2023); Dong et al. (2024a) and debiasing techniques Gallegos et al. (2024); Ranaldi et al. (2024), but most of them are on English. There are a few works on multilingual settings Zhao et al. (2024a); Vashishtha et al. (2023), but such efforts are not common for Bangla. The most preliminary work on Bangla bias detection is found in the works of Sadhu et al. (2024), that includes static and contextual embeddings. Effectiveness of varied probing techniques for extracting cultural variations in pretrained LMs has been discussed in Arora et al. (2023).
3 Linguistic Characteristics of Bangla Pronouns
Unlike English and similar languages, Bangla lacks gender-specific pronouns (e.g., he, she). Instead, Bangla employs common pronouns that are used interchangeably for both male and female genders in both singular and plural forms. Moreover, the structure of Bangla sentences does not change in terms of verbs or other grammatical elements to indicate the gender of the subject, as is the case in languages like Hindi or Spanish. As a result, sentences in Bangla that do not include gender-specific nouns or proper names are inherently gender-neutral.
@@ 4 Data
We use two strategies for LLM probing: Template Based and Naturally Sourced. The template-based approach uses curated templates for gendered persona or religious group predictions for bias evaluation. Naturally sourced sentences, on the other hand, are used to make explicit predictions about groups or genders, helping to understand the LLM’s ability to interpret natural scenarios. We explain the two techniques as follows:
Template Based: We create semantically bleached templates with placeholders for specific traits, filled with adjective words from categories like Personality, Outlook, Communal, and Occupation (see Figures 6 and 9 in appendix). The adjective categories and words were validated by native Bangla-speaking authors. To explore the effect of occupation on role prediction, we intermix professions with traits in the templates. Examples in the Placeholder column of Figure 9 illustrate the process. Care was taken to avoid stereotypes, ensuring all adjectives and occupations were equally probable for any gender or religious community. For gender detection, the templates employed gender-neutral pronouns of Bangla, along with simple and context-independent sentences to obscure any clues about the gender of the person being referred to. Similarly, for detecting bias related to religious communities, the templates used common, non-specific pronouns (e.g., they/them) and avoided any contextual or identifying details that could hint at the religious affiliation of the individual mentioned in the prompt. In total, we have 2772 template sentences by combining both the categories (see Appendix 4 for detailed statistics).
Figure 2: Workflow of Filtering Naturally Sourced Data using LLM and Prompt Preparation
Naturally Sourced: The workflow of preparing the corpus for naturally sourced sentences is illustrated in Figure 1. We use the BIBED dataset Das et al. (2023), specifically the Explicit Bias Evaluation (EBE) data for naturally occurring scenarios. The sentences are structured in pairs, each containing one identifying subject from a group of either male-female words (for gender) or Hindu-Muslim words (for religion). Figure 7 (in the appendix) illustrates how sentences are grouped into ’Gender’ and ’Religion’ biases. It provides original (root) sentences, paired sentences with altered gender or religion entities, and the modifications necessary to transform them into data points.
An important limitation of the BIBED dataset is that many sentences are not equally probable for both contrasting identities due to issues such as contradictory historical facts, entity-specific information not applicable to the other, incorrect identification of gender or religion entity in the root sentences, or lack of moderation. Examples of these non-applicable scenarios are shown in Figure 8 (in Appendix). To address this, we manually curated sentences to ensure equal applicability to both identities (see Appendix C for details). Each selected root sentence was transformed into a data point by removing the main identifying subject (male-female for gender or Hindu-Muslim for religion) and converting it into a bias detection prompt. Examples of the final prompt format are provided in the Modification column of Figure 7. The prompt creation workflow is illustrated in Figure 2. After curation, 2416 pairs were retained for gender and 1535 for religion.
@@ 5.1 Model Selection
For our experiment we provide results for four state-of-the-art LLMs: Llama3-8b (version: Meta-Llama-3-8B-Instruct 333meta-llama/Meta-Llama-3-8B-Instruct) (AI@Meta, 2024), GPT-3.5-Turbo 444gpt-3-5-turbo, GPT-4o 555gpt-4o and Claude-3.5-Sonnet666anthropic/claude-3.5-sonnet. To reduce randomness, we set the temperature very low ($temp=0.1$) and restrict the maximum response length to 128. Since Bangla is a low resource language, not many models could generate the expected response we required. Some of the open source models that we used but failed to get presentable results are mentioned in the limitations section.
@@ 5.2 Prompt
In the case of template based probing, we prompt the model for gendered role or religious identity selection, and in the case of naturally sourced probing, we use fill in the blanks approach.
Template Probing: As shown in Table 5 (appendix F), LLMs are instructed to respond with a gender or religion assuming role of a Bengali person for template based probing. Each input contains a sentence with gender neutral pronoun along with one of the trait words listed in Figure 6. Input sentence templates with placeholders are explained in Figure 9.
Naturally Sourced Probing: LLMs are instructed to fill in the blank with a gender (male-female) or religion (Hindu-Muslim) reflecting the context of the input. Modification of EBE datapoints for prompt creation is shown in Figure 7.
In table 1, we provide the number of unique prompts for each model.
@@
During evaluation, the options (gender or religion prediction) provided to LLMs inside a prompt are randomly shuffled for both gender and religious entities to avoid selection bias Zheng et al. (2024).
5.3 Evaluation Metric
We employ the widely used fairness metric, Disparate Impact (DI) Feldman et al. (2015), calculated as $\frac{P(Y=1|S\neq 1)}{P(Y=1|S=1)}$. For our binary identifiers (e.g., male-female, Hindu-Muslim), DI can be applied through empirical estimation. In task Q, for category a with outcomes x and y, DI is calculated by the following formula:
@@ 5.3 Evaluation Metric
We use occurrence frequency instead of probability Zhao et al. (2024b) and adjust the metric to adjust equal proportionality in bias scores (further justification and detail is provided in appendix B):
@@
5.4 Metric Interpretation and Bias Direction
To better understand the bias score from numerical values, we provide an interpretation framework in Table 2. Greater deviation from the neutral line denotes the presence of greater bias in either directions.
Bias Type Bias Score
Positive Negative
Gender Female-biased Male-biased
Religion Hindu-biased Muslim-biased
Table 2: Interpretation of Bias Scores for Gender and Religion
@@ 6 Results and Evaluation
Figure 3: Bias Scores in role selection for multiple LLMs in the case of template based probing for gender and religion data. Positive and negative traits results are shown separately. The neutral line $(Bias\>Score=0)$ is highlighted in all the figures. The positive bias scores in figures 3(a) and 3(b) represents Female biased and in figures 3(c) and 3(d) represents Hindu biased. (Note that the results for Occupation are the same for positive and negative traits and only included in contrasting graphs for the ease of comprehending the effect of inter-mixing with other traits.)
@@ 6.1 Template Based Probing Results
Gender Bias: Our findings (Figure 3(a), 3(b)) show that GPT-3.5-Turbo is consistently biased toward females, while Llama-3 and Claude-3.5-Sonnet are biased toward males across both positive and negative traits. GPT-4o exhibits the most fluctuation, switching its bias depending on the category. When the traits change from positive to negative, GPT-4o changes substantially from female direction to male direction for Personality and Communal based traits. Except for GPT-3.5-Turbo, all models display a strong male bias for occupations.
@@ 6.2 Naturally Sourced Probing Results
Figure 4: Bias results in Naturally Sourced(EBE) probing method for multiple LLMs
@@
We hypothesize that, the reason for not showing substantial bias in naturally probed examples can be attributed to two points: (1) When a Bangla prompt is provided with a broader and naturally occurring context, the LLMs tend to focus on the overall meaning of the scenario rather than isolating specific characters and attributing gender or religious identities to them. This reduces the likelihood of bias being explicitly reflected in the responses. (2) The guard-rails used in LLMs work better in a natural probing setting.
Key Take-away: The study reveals significant biases in multilingual large language models (LLMs) when generating outputs in Bangla. Gender and religious biases are evident, varying in degree and direction depending on the model and probing method. Template-based probing shows more pronounced biases as opposed to naturally sourced probing.
@@ 7 Conclusion
To summarize, our study investigates gender and religious bias in multilingual LLMs within the context of Bangla, utilizing two distinct probing techniques and datasets. The results reveal varying degrees of bias across models and underscore the need for effective debiasing techniques to ensure the ethical use of LLMs in sensitive Bangla-language applications. Additionally, the findings highlight the importance of developing linguistically and culturally aware frameworks for bias measurement. Future research could focus on expanding the dataset to include non-binary genders, additional religious groups, and nuanced sociocultural contexts to better capture the diversity of Bangla-speaking regions.
@@ Appendix B Evaluation Metric Justification
Appendix C Data Filtration for Naturally Sourced Sentences
The selection criteria for the Explicit Bias Evaluation(EBE) dataset are based on ensuring meaningful and contextually accurate sentences that are neutral from the perspective of gender and religion. In the original BIBED dataset Das et al. (2023), authors created pair for each sentence by replacing the identifying subject, either male-female (for gender) or Hindu-Muslim (for religion) with their respective counterparts (shown in Figure 7). However, in the EBE data, there are many generated pair sentences that are semantically inconsistent for the pair subject as illustrated in the first two columns of Figure 8.
@@ Appendix C Data Filtration for Naturally Sourced Sentences
Appendix D Annotator’s Agreement on Naturally Selected Data
@@
Appendix E Dataset Statistics
@@ Appendix F Prompt Template
Figure 6: Categories of Adjective words used for templates
Figure 8: Examples of Rejected Sentence and Reason for Rejection